2 0 1 23 4 5 6 7 8 9 10 Good AYour budget line for goods A and B is represented below. Which of the following statements is or could be TRUE?
(a) If income is $45, the price of Good A must always be $5.
(b) The price of Good A increases after the consumer buys 2 units.
(c) When you have 5 units of Good A, you are able to buy an additional unit of Good A if you give up 1/3 units of Good B.
(d) When you have 5 units of good B, the marginal rate of transformation of A for B is 1/3 (???ab = 1/3)
(e) All of the above statements are false.
Answer #1

Ans. a) If income is $45, the price of goods A must always be $5

when income is $45, the consumer can consume a maximum of 9 units of goods A

then, Qd = M/P, when income is $45= $45/9 = $5